OpenStudy (anonymous):

Help Please! Perform the indicated operation 5 1/6 - 2 2/3 I did the math (in comment) but it doesn't seem to be right

OpenStudy (anonymous):

\[\frac{ 1*3 }{6*3 } - \frac{ 2*6 }{3*6 } = \frac{ 3 }{ 18 } - \frac{ 6 }{ 18 } \] I get 3 3/18 but the answers have 2's. I need help :(

hartnn (hartnn):

where did 5 and 2 go in your calculation?

hartnn (hartnn):

u cannot just do 5-2=3

OpenStudy (anonymous):

Oh? how would I add them into the equation then?

hartnn (hartnn):

ok, \[5\frac{1}{6}=\frac{5*6+1}{6}=\frac{31}{6}\] can u do the same for 2 2/3 ??

OpenStudy (anonymous):

So I would add the 2 in with the equation like you did? But, how does a two at the end come into play?

OpenStudy (anonymous):

Evaluate the following:\[\left(5+\frac{1}{6}\right)-\left(2+\frac{2}{3}\right) \]

OpenStudy (anonymous):

i got 12/3 for the 2 2/3 one

hartnn (hartnn):

wait,i would tell that my method was lenghty,and u can do 5-3=2 as far as your work in 1st comment,the only mistake u did was 6*2=6 its actually 12

OpenStudy (anonymous):

im lost now

hartnn (hartnn):

so whats \[\frac{3-12}{18}\]

OpenStudy (anonymous):

-9/18

hartnn (hartnn):

\[\frac{ 1*3 }{6*3 } - \frac{ 2*6 }{3*6 } = \frac{ 3 }{ 18 } - \frac{ 12}{ 18 } =\frac{-9}{18}=\frac{-1}{2}\] got this?

OpenStudy (anonymous):

\[2+\frac{2}{3}=\frac{2*3}{3}+\frac{2}{3}=\frac{2*3+2}{3}=\frac{8}{3} \]

OpenStudy (anonymous):

Ok, then where does the big two come from?

OpenStudy (anonymous):

\[a+\frac{b}{c}=\frac{a c+b}{c} \]

OpenStudy (anonymous):

2 1/3. 2 1/2. 2 1/6. those are the options

hartnn (hartnn):

ok,5-2=3 so your final answer would be 3 -1/2 which is same as (6-1)/2 =5/2 =2 1/2

OpenStudy (anonymous):

Oh, so the negative is what makes the 3 a 2?

OpenStudy (anonymous):

\[\left(5+\frac{1}{6}\right)-\left(2+\frac{2}{3}\right)=\left(\frac{31}{6}\right)-\left(\frac{8}{3}\right)=\left(\frac{31}{6}\right)-\left(\frac{16}{6}\right)=\frac{15}{6}=\frac{5*3}{2*3}=\frac{5}{2} \]

OpenStudy (phi):

Do you know how to change \[ 2 \frac{2}{3} \] to an improper fraction? 2 can be changed to \(\frac{6}{3} \) so 2 +2/3 is the same as 6/3 + 2/3= 8/3